The following table shows the demand curve facing a monopolist who produces at a constant marginal cost of $8.00. Show all work.

TR MR Price Quantity 180 16 14 8 12 12 10 16 20 6 24 4 28 2 32 0 36 8

What is the firm's profit? Profit equals $____________ What would the equilibrium price and quantity be in a competitive industry?
Answer #1
Price Quantity TR MR MC ATC TC Profit
18 0 0
16 4 64 16 8 8 32 32
14 8 112 12 8 8 64 48
12 12 144 8 8 8 96 48
10 16 160 4 8 8 128 32
8 20 160 0 8 8 160 0
6 24 144 -4 8 8 192 -48
4 28 112 -8 8 8 224 -112
2 32 64 -12 8 8 256 -192
0 36 0 -16 8 8 288 -288

here the MC is constant so it will be equal to ATC.

What is the firm's profit? Profit equals $48.

explanation:

profit=TR-TC

=144-(8*12)

=144-96

=48.

What would the equilibrium price and quantity be in a competitive industry?

Price=12

quantity=12

Explanation:

firm maximizes its profit where MR=MC.

so at quantity and price 12, firm maximizes its profit as MR=MC.